solve using the elimination method. 7x-6y= -88.5 and -28x+24y= -354. I'm so lost as to how to do these, can someone please help??
Answer by AnlytcPhil(1806)

  7x -  6y =  -88.5  
-28x + 24y = -354

Multiply the first equation through by 4 to make the x's cancel

 28x - 24y = -354  
-28x + 24y = -354
-----------------
  0x +  0y = -708
         0 = -708

That is false so there is no solution.  This is an inconsistent
system of equations.

Edwin